OpenStudy (anonymous):

absolute maximum and minimum for 5x^2-6x+5 on interval [0,5]

zepdrix (zepdrix):

What part are you getting stuck on Jose? We can find max/min points when \(\large f'(x)=0\). Solving for x will get us the max/min points that we need. Have you taken the derivative of the function?

zepdrix (zepdrix):

Since it's a closed interval, I guess we'll also have to check the end points. We'll deal with that after the derivative portion though :)

OpenStudy (anonymous):

yes i have derivative is 10x-6 and i got x=6/10 i do not know where to go from here

zepdrix (zepdrix):

So we found 1 critical point. So we'll have three total points to check. The critical point that we found, and the boundary values. We'll plug each of these into the `original` function, and see which gives us the largest output, and which the smallest. It's as simple as that!

zepdrix (zepdrix):

\[\large f(x)=5x^2-6x+5\] \[\large f\left(\frac{6}{10}\right)=?\]\[\large f(0)=?\]\[\large f(5)=?\]

OpenStudy (anonymous):

F(6/10)=23/10 f(0)=5 f(5)=100

zepdrix (zepdrix):

Hmm so it looks like x=6/10 gave you the smallest value ~ your minimum. And x=5 gave you the largest ~ your maximum. It's probably good to write your answer in ordered pair notation. \(\large min=(6/10,\quad 23/10)\) \(\large max=(5,\quad 100)\)

OpenStudy (anonymous):

yea that was what I had thought, it was because this online site for my homework asked for values so i had been putting in x values, yet wen i put solely the y value it accepted the answer. thanks for your help.

zepdrix (zepdrix):

cool c: